OpenStudy (anonymous):

sqrt(x) - 1 * sqrt(x) + 1 = x right

OpenStudy (anonymous):

x - 1 *

OpenStudy (aravindg):

yes

OpenStudy (aravindg):

(a+b)(a-b)=(a^2-b^2)

OpenStudy (anonymous):

x*sqrt(x) + 1/x -1 = -2 right with limit x aproachign 1